# Sense4FL: Vehicular Crowdsensing Enhanced Federated Learning for Autonomous Driving

### v1

@@ VI-B An Approximate Algorithm
$Obj=D_{\text{local}}+D_{\text{global}},$ (20)
@@
$D_{\text{local}}=\delta\sum_{v=1}^{V}\frac{a_{v}\rho_{v}}{\sum_{v=1}^{V}a_{v}\rho_{v}}{D_{v,\text{local}}}.$ (21)
We observe that $D_{\text{local}}$ is a separable term of decision variables because it is the summation of the local EMD divergence of selected vehicles. In contrast, $D_{\text{global}}$ is non-separable, which is more challenging to optimize. Moreover, as both $D_{\text{local}}$ and $D_{\text{global}}$ quantify EMD divergence, we observe that the first term is at least a constant ratio of the second term, and will dominate when $\delta$ or local update step $T$ is large. This inspires us to develop a two-step optimization algorithm: 1) minimize the separable term $D_{\text{local}}$ first, for which we can achieve the minimum value of $D_{\text{local}}$ due to the tractability of the expression, and 2) use a local search procedure to improve the algorithm by evaluating $D_{\text{local}}+D_{\text{global}}$ until no improvement can be made. In this way, a good upper bound can be achieved for the minimization problem.
@@ VI-B1 Step 1
By focusing on $D_{\text{local}}$, the optimization problem becomes
$\min_{\mathbf{a},\mathbf{g}_{v}}~{}$ $D_{\text{local}}$ (22a)
@@
We can get rid of $\mathbf{g}_{v}$ because the optimal value $\mathbf{g}_{v}^{\star}$ can be obtained by traversing the feasible set for each vehicle independently with a low time complexity $\mathcal{O}(\sum_{v=1}^{V}\sum_{m=1}^{M_{v}}N_{v,m})$. The resultant problem becomes
@@
where $\tilde{d}_{v}=\sum_{m=1}^{M_{v}}\overline{\xi_{v,m}}\sum_{i=1}^{C}\|p_{v,m}^{i}-\sum_{b=1}^{B}l_{b}p_{b}^{i}\|$ is the local divergence with the optimal data collection scheme $\mathbf{g}_{v}^{\star}$. We observe that the problem is an integer fractional programming. To address this rather hard problem, we introduce an auxiliary variable $d$ and decompose Problem (23) into two subproblems. Specifically, by introducing $d$, the optimization problem can be reformulated as follows
@@
The second subproblem is to find the minimum value of $d$, denoted by $d^{\dagger}$, under which there is a feasible solution to (25). Since $\tilde{d}_{v}\in[0,2],1\leq v\leq{V}$, we have $d\in[0,2\delta]$. To compute $d^{\dagger}$, we use the bisection method over $[0,2\delta]$, with an error tolerance of $\sigma$. Given $d^{\dagger}$, vehicles are sorted in ascending order of $\rho_{v}(\delta\tilde{d}_{v}-d^{\dagger})$, and the first $S$ vehicles are selected. This selection strategy yields the optimal solution to Problem (23), with $d^{\dagger}$ being the optimal objective value. We have the following Theorem. .
@@
Input: $\mathcal{V}$, $\mathcal{T}$, $T^{\mathrm{task}}$, $S$, $p_{b}^{i}$, $l_{b}$, $\mathcal{H}_{v}$, $N_{\mathrm{ite}},\delta,\sigma$
@@ 5       Optimize $\mathbf{g}_{v}$ to minimize $\rho_{v}(\delta\tilde{d}_{v}-d)$;
7       Calculate $D_{\text{local}}^{\dagger}\leftarrow\sum_{v=1}^{S}\rho_{v}(\delta\tilde{d}_{v}-d)$;
8       if $D_{\text{local}}^{\dagger}\leq 0$ then
9             $u_{r}\leftarrow u$;
@@ 27                   $\mathcal{S}^{\prime}\leftarrow\mathcal{S}_{0}\setminus\{v\}\cup\{c\}$;
28                   Optimize $\mathbf{g}_{c}$ by traversing the feasible set;
@@ VI-B2 Step 2
Since Step 1 only considers the term $D_{\text{local}}$, in what follows, we refine the algorithm by considering the original objective function $D_{\text{local}}+D_{\text{global}}$ based on a local search procedure. Specifically, during each iteration, we replace the least effective vehicle, i.e., introducing the maximum incremental value to the objective function $D_{\text{local}}+D_{\text{global}}$, with the one that has the minimum value in the unselected vehicle set. This procedure continues until no further improvements can be made or the maximum number of iterations is reached. The proposed two-step algorithm is presented in Algorithm 1.
@@
We establish the following relationship between $D_{\text{local}}$ and $D_{\text{global}}$.
@@
$\leq$ $\frac{1}{\delta}D_{\text{local}}.$
From Step 1, the optimal value of $D_{\text{local}}$ is $D_{\text{local}}^{\dagger}=d^{\dagger}$, and in this case, we can calculate the value of $D_{\text{global}}^{\dagger}$ based on the corresponding solution $\mathbf{a}^{\dagger}$. Moreover, we assume the true optimal value for the original problem is $D_{\text{local}}^{\star}+D_{\text{global}}^{\star}$, which is obtained by jointly optimizing both terms. It holds that $D_{\text{local}}^{\dagger}\leq D_{\text{local}}^{\star}$. Since $D_{\text{global}}\leq\frac{1}{\delta}D_{\text{local}}$ as derived in (27), we can get $D_{\text{global}}^{\dagger}\leq\frac{1}{\delta}D_{\text{local}}^{\dagger}$. Therefore, we arrive at
$\frac{D_{\text{local}}^{\dagger}+D_{\text{global}}^{\dagger}}{D_{\text{local}}^{\star}+D_{\text{global}}^{\star}}\leq\frac{D_{\text{local}}^{\dagger}(1+\frac{1}{\delta})}{D_{\text{local}}^{\star}+D_{\text{global}}^{\star}}$ (28)
$\leq$ $\frac{D_{\text{local}}^{\dagger}(1+\frac{1}{\delta})}{D_{\text{local}}^{\star}}\leq\frac{D_{\text{local}}^{\star}(1+\frac{1}{\delta})}{D_{\text{local}}^{\star}}$
@@
Figure 3: Experiment results for training from scratch based on images captured in Singapore’s One-North in the nuImages dataset.
VII Experiments
In this section, we provide numerical experiments to evaluate the performance of our proposed Sense4FL framework. We compare our Sense4FL framework with three benchmark methods and demonstrate the superiority of our scheme.
VII-A Experiment Settings
We conduct experiments with the nuImages dataset, an extension of nuScenes designed for 2D object detection [53]. The images were captured by six cameras mounted around a vehicle, and data was collected in Singapore and Boston, respectively. Each image’s bounding box was converted to YOLO format, where a bounding box is represented by normalized center coordinates, box width, and height [41]. We divide the images based on their collection locations and generate vehicle traffic using the SUMO simulator. Moreover, the images captured in Singapore’s One-North area are designed for training from scratch. After that, we fine-tune the model in Boston scenario to illustrate the necessity for fine-tuning models based on Sense4FL when the environment has been changed. Specifically, images captured in Singapore’s One-North area are divided into 36 blocks. From these blocks, we generate 10 possible trajectories. Each vehicle has two possible trajectories, which can be predicted based on their current locations and orientations with historical traffic data. We repeat our simulation process 15 times to obtain the average performance. Besides, unless specified otherwise, the default settings are provided as follows. The total number of available vehicles in each round is around 100, and the number of vehicles to be selected is 10. The computing capability of each vehicle is 10 GFLOPS [54]. The number of parameters of the adopted YOLOv7 is 36.9 million and we adopt the 16-bit version [55]. The time constraint of the FL task for one round is 30 seconds. Vehicles travel at speeds between 12 m/s and 20 m/s. The minimum expected data rate for uploading is 40 Mbps. Each selected vehicle performs $2$-steps local SGD updates before uploading, with the batch size set to 32. The learning rate is set to 0.001.
@@ VII-A Experiment Settings
Figure 5: A fine-tuning process based on the nuImages dataset. We adopt a pre-trained model from images captured in Singapore’s One-North and then fine-tune it in Boston scenario.
@@ VII-B Performance Evaluation of Sense4FL Framework
To validate the effectiveness of the proposed Sense4FL framework, we compare it against three benchmark methods.
@@
  • Random Selection: In this method, we randomly select $S$ vehicles and their data collection in each round from all available vehicles to participate in FL training.
Fig. 3 illustrates the training loss and accuracy under different cases. The training loss is derived from the average of the local losses computed by each vehicle selected to participate in the FL process. The mean average precision (mAP) is measured by evaluating the global model at the end of each round on a separate validation dataset stored on the FL server. As seen from Fig. 3 (a)-(c), the training performance of our Sense4FL scheme considerably outperforms other benchmarks. This is because our scheme largely overcomes the model bias resulting from the inappropriate selection of vehicles and their training data and, therefore, makes the object detection model better represent the region of interest. Fig. 3 (d)-(f) evaluate the performance by varying the network settings, including the number of available vehicles, the uploading data rate, and the computing capabilities. Intuitively, where there are more vehicles available, there exists more flexibility to select better vehicles with desired training data, thereby enhancing the object detection performance. Similarly, higher uploading data rates or more powerful onboard computing capabilities increase the probability of successful reception of models by reducing the communication-computing latency in each training round, thus leading to better training performance. We also find that the proposed Sense4FL outperforms the other three baselines since the training data distribution and the uploading probabilities are both taken into account in vehicle selection. These baselines perform worse as they introduce model bias caused by the non-representative training datasets collected by vehicles.
@@
Furthermore, to demonstrate the necessity and effectiveness of our Sense4FL framework, i.e., incorporating trajectory-dependent training data collection and vehicle selection for FL to improve AD performance, we fine-tune the model pre-trained from Singapore’s One-North area into another region, i.e., Boston. As shown in Fig. 5, the phenomenon is similar to what we can observe from Fig. 3. Specifically, the model performance of our scheme substantially outperforms other benchmarks in the fine-tuning process. We can take three insights from Fig. 5. First, the AD model well-trained for one area may not perform very well under the other scenario, implying that a general AI model is not applicable to all cases and fine-tuning is needed. Second, the time needed for convergence is much shorter. This highlights the feasibility of fine-tuning an object detection model to achieve satisfactory performance in a new environment within a reasonable time frame. At last, the proposed Sense4FL is a trajectory-dependent approach, which outperforms other methods, as it overcomes the model bias resulting from the inappropriate selection of vehicles.

### v2

@@ VI-B An Approximate Algorithm
$Obj=D_{\text{client}}+D_{\text{global}},$ (21)
@@
$D_{\text{client}}=\delta\sum_{v=1}^{V}\frac{a_{v}\rho_{v}}{\sum_{v=1}^{V}a_{v}\rho_{v}}{D_{v,\text{client}}}.$ (22)
We observe that $D_{\text{client}}$ is a separable term of decision variables because it is the summation of the local EMD divergence of selected vehicles. In contrast, $D_{\text{global}}$ is non-separable, which is more challenging to optimize. Moreover, as both $D_{\text{client}}$ and $D_{\text{global}}$ quantify EMD divergence, we observe that the first term is at least a constant ratio of the second term, and will dominate when $\delta$ or local update step $T$ is large. This inspires us to develop a two-step optimization algorithm: 1) minimize the separable term $D_{\text{client}}$ first, for which we can achieve the minimum value of $D_{\text{client}}$ due to the tractability of the expression, and 2) use a local search procedure to improve the algorithm by evaluating $D_{\text{client}}+D_{\text{global}}$ until no improvement can be made. In this way, a good upper bound can be achieved for the minimization problem.
@@ VI-B1 Step 1
By focusing on $D_{\text{client}}$, the optimization problem becomes
$\min_{\mathbf{a},\mathbf{g}_{v}}$~ $D_{\text{client}}$ (23a)
@@
We can get rid of $\mathbf{g}_{v}$ because the optimal value $\mathbf{g}_{v}^{\star}$ can be obtained by traversing the feasible set for each vehicle independently with a low time complexity $\mathcal{O}(\sum_{m=1}^{M_{v}}N_{v,m})$. The resultant problem becomes
@@
where $\tilde{d}_{v}=\sum_{m=1}^{M_{v}}\overline{\xi_{v,m}}\sum_{i=1}^{C}\|p_{v,m}^{i}-\sum_{b=1}^{B}l_{b}p_{b}^{i}\|$ is the client divergence with the optimal data collection scheme $\mathbf{g}_{v}^{\star}$. We observe that the problem is an integer fractional programming. To address this rather hard problem, we introduce an auxiliary variable $d$ and decompose Problem (24) into two subproblems. Specifically, by introducing $d$, the optimization problem can be reformulated as follows
@@
The second subproblem is to find the minimum value of $d$, denoted by $d^{\dagger}$, under which there is a feasible solution to (26). Since $\tilde{d}_{v}\in[0,2],1\leq v\leq{V}$, we have $d\in[0,2\delta]$. To compute $d^{\dagger}$, we use the bisection method over $0,2\delta]$, with an error tolerance of $\sigma$. Given $d^{\dagger}$, vehicles are sorted in ascending order of $\rho_{v}(\delta\tilde{d}_{v}-d^{\dagger})$, and the first $S$ vehicles are selected. This selection strategy yields the optimal solution to Problem (24), with $d^{\dagger}$ being the optimal objective value. We have the following Theorem.
@@
Input: $\mathcal{V}$, $\mathcal{T}$, $T^{\mathrm{task}}$, $S$, $p_{b}^{i}$, $l_{b}$, $\mathcal{H}_{v}$, $N_{\max},\delta,\sigma$
@@ 5   Optimize $\mathbf{g}_{v}$ to minimize $\rho_{v}(\delta\tilde{d}_{v}-d)$;
7   Calculate $D_{\text{client}}^{\dagger}\leftarrow\sum_{v=1}^{S}\rho_{v}(\delta\tilde{d}_{v}-d)$;
8 if $D_{\text{client}}^{\dagger}\leq 0$ then
9    $d_{r}\leftarrow d$;
@@ 26       $\mathcal{S}^{\prime}\leftarrow\mathcal{S}_{0}\setminus\{v\}\cup\{c\}$;
27         Optimize $\mathbf{g}_{v}$ by traversing the feasible set;
@@ VI-B2 Step 2
Since Step 1 only considers the term $D_{\text{client}}$, in what follows, we refine the algorithm by considering the original objective function $D_{\text{client}}+D_{\text{global}}$ based on a local search procedure. Specifically, during each iteration, we replace the least effective vehicle, i.e., introducing the maximum incremental value to the objective function $D_{\text{client}}+D_{\text{global}}$, with the one that has the minimum value in the unselected vehicle set. This procedure continues until no further improvements can be made or the maximum number of iterations is reached. The proposed two-step algorithm is presented in Algorithm 1.
@@
We establish the following relationship between $D_{\text{client}}$ and $D_{\text{global}}$.
@@
$\leq$ $\frac{1}{\delta}D_{\text{client}}.$
From Step 1, the optimal value of $D_{\text{client}}$ is $D_{\text{client}}^{\dagger}=d^{\dagger}$, and in this case, we can calculate the value of $D_{\text{global}}^{\dagger}$ based on the corresponding solution $\mathbf{a}^{\dagger}$. Moreover, we assume the true optimal value for the original problem is $D_{\text{client}}^{\star}+D_{\text{global}}^{\star}$, which is obtained by jointly optimizing both terms. It holds that $D_{\text{client}}^{\dagger}\leq D_{\text{client}}^{\star}$. Since $D_{\text{global}}\leq\frac{1}{\delta}D_{\text{client}}$ as derived in (28), we can get $D_{\text{global}}^{\dagger}\leq\frac{1}{\delta}D_{\text{client}}^{\dagger}$. Therefore, we arrive at
$\frac{D_{\text{client}}^{\dagger}+D_{\text{global}}^{\dagger}}{D_{\text{client}}^{\star}+D_{\text{global}}^{\star}}\leq\frac{D_{\text{client}}^{\dagger}(1+\frac{1}{\delta})}{D_{\text{client}}^{\star}+D_{\text{global}}^{\star}}$ (29)
$\leq$ $\frac{D_{\text{client}}^{\dagger}(1+\frac{1}{\delta})}{D_{\text{client}}^{\star}}\leq\frac{D_{\text{client}}^{\star}(1+\frac{1}{\delta})}{D_{\text{client}}^{\star}}$
@@
Finally, we analyze the computational complexity of the proposed algorithm.
Theorem 5.
The overall computational complexity for Algorithm 1 is $\mathcal{O}\left((VN_{\mathrm{tot}}\log\frac{2V}{\sigma}+N_{\mathrm{ite}}N_{\mathrm{tot}}S(V-S))\right)$, where $N_{\mathrm{tot}}=\sum_{v=1}^{V}\sum_{m=1}^{M_{v}}N_{v,m}$.
Proof.
We propose a two-step algorithm to optimize both vehicle selection and data collection strategy. In Step 1, a bisection method is employed to find the optimal vehicle selection. This process has a computational complexity of $\mathcal{O}\left(VN_{\mathrm{tot}}\log\frac{2V}{\sigma}\right)$, where $N_{\mathrm{tot}}=\sum_{v=1}^{V}\sum_{m=1}^{M_{v}}N_{v,m}$. To refine the selection by considering the original objective function, we introduce a local search procedure, with a complexity of $\mathcal{O}\left(N_{\mathrm{ite}}N_{\mathrm{tot}}S(V-S)\right)$, where $N_{\mathrm{ite}}$ is the number of iterations until convergence. Consequently, the overall computational complexity for Algorithm 1 is $\mathcal{O}\left((VN_{\mathrm{tot}}\log\frac{2V}{\sigma}+N_{\mathrm{ite}}N_{\mathrm{tot}}S(V-S))\right)$. ∎
The low complexity of the algorithm, combined with the parallel processing of vehicles, ensures the scalability of our proposed framework even in dense urban scenarios.
TABLE III: Parameter settings for simulations.
Number of selected vehicles $S$ = 10
Model size $\omega=5.904\times 10^{8}$ bit
Number of processing cycles $c_{v}=9.8304\times 10^{7}$
Vehicle speed 40-50 km/h / 50-60 km/h
Time constraint for one round $T^{\text{task}}=80$ seconds
Number of possible trajectories $M_{v}=2$
Steps of local SGD updates $T=2$
Time required via wired link $t^{\mathrm{trans}}$ = 1 second
Computing capability $f_{v}$ = 40 GFLOPS
Minimum uploading data rate 50 Mbps
Lipschitz parameter $\lambda_{\text{max}}=0.01$
Batch size $D_{\text{Batch}}=32$
VII Experiments
In this section, we provide numerical experiments to evaluate the performance of our proposed Sense4FL framework. We compare our Sense4FL framework with several benchmark methods and demonstrate the superiority of our scheme.
VII-A Experiment Settings
We conduct experiments with the nuImages dataset, an extension of nuScenes designed for 2D object detection [48]. The images were captured by six cameras mounted around a vehicle, and data was collected in Singapore and Boston, respectively. Each image’s bounding box was converted to YOLO format, where a bounding box is represented by normalized center coordinates, box width, and height [39]. We divide the images based on their collection locations and generate vehicle traffic using the SUMO simulator. Moreover, the images captured in Singapore’s One-North area are designed for training from scratch. After that, we adapt the model in the Boston Seaport scenario to illustrate the necessity for adapting models based on Sense4FL when the environment has changed. Specifically, the data is split into training/adaptation and test sets with a ratio of 4:1. Images are divided into 36 distinct blocks based on their capture location. From these blocks, we generate 10 realistic trajectories, which are illustrated in Fig. 4. This location-dependent data splitting strategy naturally induces heterogeneity, which can be observed for each trajectory in both Singapore’s One-North and Boston Seaport in Fig. 5. We repeat our simulation process 15 times to obtain the average performance.
(a) The trajectories in Singapore’s One-North.
(b) The trajectories in Boston Seaport.
Figure 4: The trajectories in Singapore’s One-North and Boston Seaport.
(a) The normalized frequency of objects for each trajectory in Singapore’s One-North.
(b) The normalized frequency of objects for each trajectory in Boston Seaport.
Figure 5: The data distribution for each trajectory in Singapore’s One-North and Boston Seaport. The data is split non-IID among ten trajectories based on the data capture location. The different colors represent different classes: pedestrian, barrier, traffic cone, bicycle, bus, car, construction vehicle, motorcycle, trailer, and truck.
@@ VII-A Experiment Settings
Figure 6: Experiment results for training from scratch in Singapore’s One-North scenario in the nuImages dataset.
Besides, unless specified otherwise, the default settings are provided as follows. The number of available vehicles is around 60-70 in Singapore’s One-North and 70-80 in Boston Seaport, and vehicles in Singapore’s One-North travel at speeds between 50-60 km/h, while those in Boston Seaport travel at 40-50 km/h. We set the maximum acceleration at 2.0 $m/s^{2}$, the maximum deceleration (braking) at 3.0 $m/s^{2}$ based on the Krauss model. Each vehicle has $M_{v}=2$ possible trajectories, which can be predicted based on its current location and orientation with historical traffic data. The number of vehicles to be selected is $S$ = 10. We consider 640*640 images with a color depth of 24 bits, and thus the required number of processing cycles for computing one sample is $c_{v}=9.8304\times 10^{7}$ [45]. The computing capability of each vehicle is $f_{v}$ = 40 GFLOPS. The adopted YOLOv7 model, with 36.9 million parameters and using a 16-bit version, has a model size of $\omega=5.904\times 10^{8}$ bits. The time constraint of the FL task for one round is $T^{\text{task}}=80$ seconds. The minimum expected data rate for uploading is 50 Mbps, and the time required for the BS to transmit the model to the FL server via a wired link is $t^{\mathrm{trans}}$ = 1 second. Each selected vehicle performs $T=2$ steps local SGD updates before uploading, with the batch size $D_{\text{Batch}}=32$. The learning rate for training from scratch is set to $\eta=0.001$ and for adaptation is set to $\eta=0.0001$. For the hyperparameter, the Lipschitz parameter $\lambda_{\text{max}}$ can be estimated and is 0.01 in this model [42]. The key parameters are summarized in Table III.
@@
Figure 8: Experiment results for model adaptation. We adapt a model pre-trained in Singapore’s One-North to the Boston Seaport scenario.
@@ VII-B Performance Evaluation of Sense4FL Framework
To validate the effectiveness of the proposed Sense4FL framework, we compare it against several benchmark methods.
  • Gradient-based selection [49]. This method selects vehicles with the highest norms of gradient values at each communication round. A vehicle stops data collection once the edge server has selected it.
  • Power-of-Choice [50]. This framework selects vehicles based on their local loss value. Upon selection by the edge server, the vehicle ceases its data collection process.
@@
  • Random Selection [9]: In this method, we randomly select $S$ vehicles and their data collection in each round from all available vehicles to participate in FL training.
  • Centralized ML: We also compare the results with the centralized machine learning (ML), which serves as the performance upper bound as the server can access all training data in this specific region.
Fig. 6 illustrates the training loss and accuracy under different cases. The training loss is derived from the average of the local losses computed by each vehicle selected to participate in the FL process. The mean average precision (mAP) is measured by evaluating the global model at the end of each round on a separate test dataset stored on the FL server. As seen from Fig. 6 (a)-(c), the training performance of our Sense4FL scheme considerably outperforms other benchmarks. This is because our scheme largely overcomes the model bias resulting from the inappropriate selection of vehicles and their training data and, therefore, makes the object detection model better represent the region of interest. Fig. 6 (d)-(f) evaluate the performance by varying the network settings, including the number of available vehicles, the uploading data rate, and the computing capabilities. Intuitively, where there are more vehicles available, there exists more flexibility to select better vehicles with desired training data, thereby enhancing the object detection performance. Similarly, higher uploading data rates or more powerful onboard computing capabilities increase the probability of successful reception of models by reducing the communication-computing latency in each training round, thus leading to better training performance. We also find that the proposed Sense4FL outperforms the other baselines since the training data distribution and the uploading probabilities are both taken into account in vehicle selection. These baselines perform worse as they introduce model bias caused by the non-representative training datasets collected by vehicles.
@@
In reality, autonomous driving companies can have a pre-trained model (say, from another city or from a cloud-based pre-training dataset) and then deploy the model in a new environment. To reflect this, we adapt the model pre-trained from Singapore’s One-North area into another region, i.e., Boston Seaport. As shown in Fig. 8, the phenomenon is similar to what we can observe from Fig. 6. Specifically, the model performance of our scheme substantially outperforms other benchmarks in the model adaptation. We can take three insights from Fig. 8. First, the AD model well-trained for one area may not perform very well under the other scenario, implying that a general AI model is not applicable to all cases and adapting is needed. Second, the time needed for convergence is much shorter. This highlights the feasibility of adapting an object detection model to achieve satisfactory performance in a new environment within a reasonable time frame. At last, the proposed Sense4FL is a trajectory-dependent approach, which outperforms other methods, as it overcomes the model bias resulting from the inappropriate selection of vehicles.
TABLE IV: Precision and Recall for different vehicle selection and data collection strategies, with the best results excluding Centralized ML shown in bold.
Singapore’s One-North Boston Seaport
Method Precision Recall Precision Recall
Proposed Sense4FL 70.16 61.28 69.13 60.27
Gradient Based 65.32 57.43 65.14 56.31
Power-of-Choice 63.29 56.17 64.28 54.24
Uploading Centric 59.86 47.81 58.23 46.18
Coverage Centric 57.77 46.40 56.57 45.64
Random Selection 52.20 41.35 50.19 40.22
Centralized ML 77.55 69.22 76.28 68.25
TABLE V: Effects of the data collection strategies, with the best results excluding Centralized ML shown in bold.
Singapore’s One-North Boston Seaport
Method mAP@0.5 mAP@0.5:.95 mAP@0.5 mAP@0.5:.95
Proposed Sense4FL 55.41 36.23 52.83 35.76
Full-data Collection 45.82 31.48 45.10 30.61
Selection Only 49.05 33.89 47.23 33.10
Centralized ML 58.38 40.12 57.42 39.76
Table IV presents the precision and recall achieved by various vehicle selection and data collection strategies. The simulation results indicate that the Sense4FL framework outperforms all baselines in avoiding false positives (high precision) and minimizing false negatives (high recall). Finally, the ablation studies on the data collection scheme in Table V reports two baselines: 1) Full-data Collection: Each vehicle collects all available data along its entire predicted trajectory, i.e., $g_{v,m}=N_{v,m}$, and we select $S$ vehicles; 2) Selection Only: This method selects $S$ vehicles but does not optimize data collection and each vehicle stops collecting data upon being selected, i.e., $g_{v,m}=c_{v,m}$. We can see that the Sense4FL framework consistently outperforms the baselines, demonstrating the salient advantage of joint optimization of vehicle selection and data collection.